OpenStudy (anonymous):

Five times the sum of a number and 27 is greater than or equal to six times the sum of that number and 26. What is the solution set of this problem?

OpenStudy (tbnrfrags):

n\[N \le -21\]

OpenStudy (tbnrfrags):

you would take \[5(N+27)\ge 6(N+26)\] and you would get \[5N+135\ge 6n+156\] then solve that and you get \[-N \ge 21\] so that would make your answer \[N \le 21\]

OpenStudy (tbnrfrags):

hopefully that helped c;

OpenStudy (anonymous):

it did